8.11 Design a full-dimensional and a reduced-dimensional state estimator for the state-space equation in Problem 8.1. Select8.1 Given X(0) = [ 21 1 c)+ [ 2 ]uco), y()= [1 1]XC)
Answer #1

ANSWER

Given thatA full-dimensional and a seduced-dimensional state Estimator for the state-space Exuation, ilt) = AxtBu, y=a let: Il be the m2x0 ) 241 +2 12-6= 8 => -e, -bo+3=8 12-3 = 16 elz= 16+3 d2=19 Substitute by value in D di+19=320 eli +12=0 l=-12 ñ = nî+Bu+L(@ [ty ta] [71- [+0] [ti tg) - 03 [1 ] (34-t2 1+ty] + (101 10t2] = [1 ] 24-t2 +10+1=” 1971 – 2 = 1 -- ti+tz+102= ti+litz=1 sol